OpenStudy (anonymous):

For each function, what is the output of the given input? For f(x) = 4x + 7, find f(4)

OpenStudy (anonymous):

20 is the answer or output

OpenStudy (anonymous):

how do you do it step by step?

OpenStudy (anonymous):

so to get F(4) you substitute 4 in the place off x. f(4)=4(4)+7 =16+7 =23

OpenStudy (anonymous):

*of